OpenStudy (anonymous):

can anyone tell me the solution of sec 11π/6 ??? FOR ADDition formulas? couz I really dont get it

Nnesha (nnesha):

\[\huge\rm sec \theta =\frac{ 1 }{ \cos \theta }\] you can use unit circle to find cos value (x,y) x= cos y=sin

Nnesha (nnesha):